Best Solutions ​Inc., sells computer hardware to end consumers. The CX30 is sold as a​ "bundle", which includes three hardware​ products: a personal computer​ (PC) tower, a​ 26-inch monitor, and a color laser printer. Each of these products is made in a separate manufacturing division of Best Solutions and can be purchased individually as well as in a bundle. Best Solutions sells roughly equal quantities of the three products. The individual selling prices and per unit costs are as​ follows:

Computer Component

Individual Selling Price Per Unit

Cost Per Unit

PC Tower

$784

$322

Monitor

$256

$182

Color Laser Print

$560

$196

Computer bundle purchase price

£900

1.

Allocate the revenue from the computer bundle purchase to each of the hardware products using the​ stand-alone method based on the individual selling price per unit.

2.

Allocate the revenue from the computer bundle purchase to each of the hardware products using the​ stand-alone method based on cost per unit.

3.

Allocate the revenue from the computer bundle purchase to each of the hardware products using the​ stand-alone method based on physical units​ (that is, the number of individual units of product sold per​ bundle).

4.

Which basis of allocation makes the most sense in this​ situation? Explain your answer.
Answer #1

1)

Items standalone price weight based on standalone price Allocation of revenue (price*weight)
PC Tower 784 784/1600= .49 900*.49= 441
Monitor 256 256/1600 = .16 900*.16= 144
Color Laser print 560 560/1600 =.35 900*.35= 315
1600 900

2)

Items standalone cost weight based on standalone cost Allocation of revenue (price*weight)
PC Tower 322 322/700=..46 900*.46=414
Monitor 182 182/700=.26 234
Color Laser print 196 196/700=.28 252
700 900

3)

Items Physical quantity weight based on standalone price Allocation of revenue (price*weight)
PC Tower 1 1/3=.33333 300
Monitor 1 1/3=.33333 300
Color Laser print 1 1/3=.33334 300

3

900

4)The Allocation based on Individual selling price per unit makes a most sensible allocation since It reflects the true % of individual selling price of products included in a Bundle or a price of bundle.